OpenStudy (anonymous):

Solve the equation: 2(3x-1)-12x-2(5-3x)=18

OpenStudy (anonymous):

You need to distribute the coefficients

OpenStudy (anonymous):

when you distribute it becomes: (6x-2)-12x-(10-6x)=18

OpenStudy (anonymous):

then you can distribute the negative signs and drop the parenthesis

OpenStudy (anonymous):

so it becomes: 6x-2-12x-10+6x=18

OpenStudy (anonymous):

then combine like-terms

OpenStudy (anonymous):

so it becomes: -12x-12=18

OpenStudy (anonymous):

add the 12 to the other side

OpenStudy (anonymous):

so it becomes: -12x=30

OpenStudy (anonymous):

divide both sides by -12 and you get x=-2.5

OpenStudy (anonymous):

Thank you very much